# Do We Need All the Synthetic Data? Towards Targeted Synthetic Image Augmentation via Diffusion Models

### v1

@@ 5.4 Ablation studies
Our method can be stacked with strong augmentation methods. Figure 3 right shows that our method stacked with TrivialAugment [29] achieves state-of-the-art results when training ResNet18 on CIFAR10. Appendix F.2 shows similar results for CIFAR100 and Tiny-ImageNet.
Effect of the upsampling factor. Figure 4(a) illustrates the performance of our method and Useful on CIFAR10 when varying the upsampling factors. As discussed in Section 3.2, Useful achieves the best performance at $k=2$ due to overfitting noise at larger $k$. On the other hand, using synthetic images, our method benefits from larger values of $k$, yielding the best performance at $k=5$ for CIFAR10 and CIFAR100, and $k=4$ for Tiny-ImageNet. Detailed results for CIFAR100 and Tiny-ImageNet can be found in Table 2 in Appendix F.2. This corroborates our theoretical findings in Section 4.2.
Choices of initialization for denoising. To validate the necessity of our faithful synthetic image generation pipeline, we run a baseline where synthetic images are generated using the standard diffusion pipeline, i.e., denoised from random noises instead of the real images. Figure 4(b) illustrates that faithful synthetic images are necessary for targeted synthetic image augmentation. When using random noises to generate synthetic images, the performance of ResNet18 on the augmented training datasets is even worse than that on the original set. This is because the generated images are not close enough to real images, resulting in a large noise which in turn harms the model performance as discussed in Theorem 4.2.
Effect of the number of denoising steps. Figure 4(c) demonstrates the effect of the number of denoising steps on the performance of our method. Using fewer steps generates images that are too close to real images, amplifying the noise in the real images. On the other hand, using too many steps results in too much noise, which also harms the performance. Overall, using 50 steps yields the best results for both SGD and SAM.
Convergence. Figure 5 in Appendix F.2 shows the mini-batch gradient variance when training ResNet18 with SGD on the augmented CIFAR10 dataset found by Useful and Ours. It can be seen that training on our augmented datasets results in lower gradient variance at every upsampling factor $k$. This confirms our theoretical results in Theorem 4.3 and Corollary 4.4.
@@ 6 Conclusion
In this work, we show that synthetically augmenting part of the data that is not learned early in training outperforms augmenting the entire dataset. By analyzing a two-layer CNN, we prove that this method allows features to be learned at a more uniform speed without amplifying noise. We conducted extensive experiments showing that our augmentation strategy boosts the performance when training ResNet, ViT and DenseNet on CIFAR-10, CIFAR-100, and TinyImageNet, with a range of optimizers including SGD and SAM, by up to 2.8%. Notably, our method applied with SGD outperforms SAM on CIFAR-100 and TinyImageNet and easily stacks with existing weak and strong data augmentation strategies to obtain state-of-the-art performance.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims match the theoretical and experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide full assumptions in both the main text and Appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide full assumptions and proofs in Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed experimental settings in the Experiment section and in Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use open datasets, which have been cited and described in the paper. We will provide our code for reproducing the experimental results in the supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide detailed experimental settings in the Experiment section and in Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the error bar for averaging multiple seeds in our experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the information on our computational resources in Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our paper conform to every aspect in the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There are no potential harms to our models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all data, code, and models used properly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not pose any potential risks to participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: We only use LLM for writing and editing purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ A.1 Full Gradient Formulas & Useful Lemmas
$\quad-\frac{3k}{N_{\text{new}}}\sum_{i=\alpha N+1}^{N}l_{i}^{(t)}\left(\beta_{d}^{3}\langle{\bm{w}}_{j}^{(t)},{\bm{v}}_{d}\rangle^{2}{\bm{v}}_{d}+y_{i}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}{\bm{\xi}}_{i}\right)$
@@
and $\text{and }\ l_{i,\bm{\epsilon}}^{(t)}=\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{d}^{3}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},{\bm{v}}_{d}\rangle^{3}-y_{i}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle^{3}\right).$
@@
With loss of generality, we consider the case when $\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle>0$. We then define the corresponding sequence $a_{t}=\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle$ generated by the update:
@@
Next, we want to show that given a proper $\eta$, this sequence is monotonic towards 00, meaning that these noises are gradually “unlearned" by the weights. We provide an inductive step below, which can be easily generalized.
@@
the weights’ alignment with these noises will be monotonic throughout the updates and get closer and closer to 00. The proof for SAM is similar (we replace $\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle$ by $\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\xi}_{i}\rangle$ for ${\bm{\xi}}_{i}\in{\mathcal{I}}_{j,\bm{\epsilon},-}^{(t)}$, etc.).
@@
And a similar expression holds for SAM. Consequently, the alignments will be monotonically away from 00, continually being learned through iterations. Hence, we want to compare how the gradients align with noises: $|\langle\nabla_{{\bm{w}}_{j,\bm{\epsilon}}}{\mathcal{L}}({\bm{W}}^{(t)}\!+\bm{\epsilon}^{(t)}),\bm{\xi}_{i}\rangle|$ vs. $|\langle\nabla_{{\bm{w}}_{j}}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\xi}_{i}\rangle|$. By Lemma A.6, we have that
@@
For $\text{For {SAM:}}\ \ \ |\langle\nabla_{{\bm{w}}_{j,\bm{\epsilon}}}{\mathcal{L}}({\bm{W}}^{(t)}\!+\bm{\epsilon}^{(t)}),\bm{\xi}_{i}\rangle|$ $=\frac{3}{N}l_{i,\bm{\epsilon}}^{(t)}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\xi}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2}$
=$\overset{(i)}{=}\frac{3}{N}l_{i,\bm{\epsilon}}^{(t)}\left(\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle-\frac{3\rho^{(t)}}{N}l_{i}^{(t)}y_{i}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\|{\bm{\xi}}_{i}\|^{2}\right)^{2}\|\bm{\xi}_{i}\|^{2}$

### v2

@@ 5.2 Ablation studies
Our method stacked with strong augmentation. Figure 4 right shows that our method stacked with TrivialAugment (Müller & Hutter, 2021) achieves state-of-the-art results when training ResNet18 on CIFAR10. Appendix F.2 shows similar results for CIFAR100 and Tiny-ImageNet.
Identifying slow-learnable features. Table 7 in Appendix F.2 shows that identifying slow-learnable features by clustering model outputs outperforms selection based on high-loss or misclassification.
For larger $k$, upsampling hurts but generation helps. Figure 5(a) illustrates the performance of our method and Upsample on CIFAR10 when varying the upsampling factors. Upsampling achieves the best performance at $k=2$ due to overfitting noise at larger $k$. But using synthetic images, our method benefits from larger values of $k$, yielding the best performance at $k=5$ for CIFAR10 and CIFAR100, and $k=4$ for Tiny-ImageNet. Detailed results for CIFAR100 and Tiny-ImageNet can be found in Table 5 in Appendix F.2. This corroborates our theoretical findings in Section 4.2.
Choices of initialization for denoising. We compare generating synthetic images by adding noise and denoising real images with denoising from random noises. Figure 5(b) illustrates that real data guidance is necessary for targeted synthetic image augmentation. When using random noises to generate synthetic images, performance of ResNet18 on the augmented training datasets is even worse than that on the original set, as generated images do not effectively amplify slow-learnable features.
Number of denoising steps. Figure 5(c) demonstrates the effect of the number of denoising steps on the performance of our method. Using fewer steps generates images that are too close to real images, amplifying the noise in the real images. In contrast, using too many steps results in too much noise, which also harms the performance. Overall, using 50 steps yields the best results for both SGD and SAM.
Convergence. Figure 7 in Appendix F.2 shows the mini-batch gradient variance when training ResNet18 on the augmented CIFAR10 dataset using upsampling and generation. Generation results in lower gradient variance compared to upsampling for different $k$, confirming our results in Sec. 4.3.
@@ 6 Conclusion
In this work, we show that synthetically augmenting part of the data that is not learned early in training outperforms augmenting the entire dataset. By analyzing a two-layer CNN, we prove that this method allows features to be learned at a more uniform speed without amplifying noise. We conducted extensive experiments showing that our augmentation strategy boosts the performance when training ResNet, ViT, DenseNet, ConNexXt and Swin Transformer on CIFAR-10, CIFAR-100, and TinyImageNet, with a range of optimizers including SGD and SAM, by up to 2.8%. Notably, our method applied with SGD outperforms SAM on CIFAR-100 and TinyImageNet and easily stacks with existing weak and strong data augmentation strategies to obtain state-of-the-art performance.
@@ A.1 Full Gradient Formulas & Useful Lemmas
$\quad-\frac{3k}{N_{\text{new}}}\sum_{i=\alpha N+1}^{N}l_{i}^{(t)}\left(\beta_{d}^{3}\langle{\bm{w}}_{j}^{(t)},{\bm{v}}_{d}\rangle^{2}{\bm{v}}_{d}+y_{i}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}{\bm{\xi}}_{i}\right),$
@@
$\text{and }\ l_{i,\bm{\epsilon}}^{(t)}=\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{d}^{3}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},{\bm{v}}_{d}\rangle^{3}-y_{i}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle^{3}\right).$
@@
Similar to Nguyen et al. (2024), we assume that the logit terms are controlled in some sense to reduce the nonlinearity. More formally, for Theorem 4.1, we will use the approximation $l_{i}^{(t)}=l_{i,\bm{\epsilon}}^{(t)}=\Theta(1)$ for any early iteration $t$. At a high level, this follows from the small perturbation and the fact that the weights have not “significantly” learned any noises (despite some overfitting) in the early phase.
$\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{d}^{3}\langle{\bm{w}}_{j}^{(t)},{\bm{v}}_{d}\rangle^{3}-y_{i}\langle{\bm{w}}_{j}^{(t)},\bm{\phi}_{i}\rangle^{3}\right)=\text{sigmoid}\left(\sum_{j=1}^{J}-\beta_{d}^{3}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},{\bm{v}}_{d}\rangle^{3}-y_{i}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle^{3}\right).$
Hence, this is implicitly assuming a small $\rho$ such that ${\bm{w}}_{j}^{(t)}$ and ${\bm{w}}_{j,\bm{\epsilon}}^{(t)}$ are not too different, and a large number of neurons $J$ such that the noise term in the sigmoid $\sum_{j=1}^{J}y_{i}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\phi}_{i}\rangle^{3}$ is ‘more ‘diluted” (close to 0).
@@
Without loss of generality, we consider the case when $\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle>0$. We then define the corresponding sequence $a_{t}=\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle$ generated by the update:
@@
Next, we want to show that given a proper $\eta$, this sequence is monotonic towards 0, meaning that these noises are gradually “unlearned” by the weights. We provide an inductive step below, which can be easily generalized.
@@
the weights’ alignment with these noises will be monotonic throughout the updates and get closer and closer to 0. The proof for SAM is similar (we replace $\langle{\bm{w}}_{j}^{(t)},\bm{\xi}_{i}\rangle$ by $\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\xi}_{i}\rangle$ for ${\bm{\xi}}_{i}\in{\mathcal{I}}_{j,\bm{\epsilon},-}^{(t)}$, etc.).
@@
And a similar expression holds for SAM. Consequently, the alignments will be monotonically away from 0, continually being learned through iterations. Hence, we want to compare how the gradients align with noises: $|\langle\nabla_{{\bm{w}}_{j,\bm{\epsilon}}}{\mathcal{L}}({\bm{W}}^{(t)}\!+\bm{\epsilon}^{(t)}),\bm{\xi}_{i}\rangle|$ vs. $|\langle\nabla_{{\bm{w}}_{j}}{\mathcal{L}}({\bm{W}}^{(t)}),\bm{\xi}_{i}\rangle|$. By Lemma A.6, we have that
@@
$\text{For {SAM:}}\ \ \ |\langle\nabla_{{\bm{w}}_{j,\bm{\epsilon}}}{\mathcal{L}}({\bm{W}}^{(t)}\!+\bm{\epsilon}^{(t)}),\bm{\xi}_{i}\rangle|$ $=\frac{3}{N}l_{i,\bm{\epsilon}}^{(t)}\langle{\bm{w}}_{j,\bm{\epsilon}}^{(t)},\bm{\xi}_{i}\rangle^{2}\|\bm{\xi}_{i}\|^{2}$
$\overset{(i)}{=}\frac{3}{N}l_{i,\bm{\epsilon}}^{(t)}\left(\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle-\frac{3\rho^{(t)}}{N}l_{i}^{(t)}y_{i}\langle{\bm{w}}_{j}^{(t)},{\bm{\xi}}_{i}\rangle^{2}\|{\bm{\xi}}_{i}\|^{2}\right)^{2}\|\bm{\xi}_{i}\|^{2}$